Calculate the number of molecules in 0.000756 grams of gaseous oxygen
1
Expert's answer
2019-10-08T08:19:48-0400

Solution.

"\frac{m}{M} = \frac{N}{Na}"

"N=\frac{m \times Na}{M}"

"N=1.42 \times 10^{19}"

Answer:

"N=1.42 \times 10^{19}"